A simple random sample of size n is drawn from a population that is normally distributed. The sample mean, is found to be 110, and the sample standard deviation S is found to be 25. Construct a 99% confidence interval for if the sample size n = 21
Answer #1

c )solution

Given that,

\bar x = 110

s =25

n =21

Degrees of freedom = df = n - 1 = 21- 1 =20

At 99% confidence level the t is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t \alpha /2,df= t0.005, 20= 2.845 ( using student t table)

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.845 * ( 25/ \sqrt 21) = 15.5207

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

110 -15.5207 < \mu < 110+ 15.5207

94.4793 < \mu < 125.1507

( 94.4793,125.1507)